# Predictive AI with External Knowledge Infusion for Stocks

### v1

@@ Abstract
Fluctuations in stock prices are influenced by a complex interplay of factors that go beyond mere historical data. These factors, themselves influenced by external forces, encompass inter-stock dynamics, broader economic factors, various government policy decisions, outbreaks of wars, etc. Furthermore, all of these factors are dynamic, and exhibit changes over time. In this paper, for the first time, we tackle the forecasting problem under external influence, by proposing learning mechanisms that not only learn from historical trends but also incorporate external knowledge from temporal knowledge graphs. Since there are no such datasets or temporal knowledge graphs available, we study this problem with stock market data, and we construct comprehensive temporal knowledge graph datasets. In our proposed approach, we model relations on external temporal knowledge graphs as events of a Hawkes process on graphs. With extensive experiments, we show that learned dynamic representations effectively rank stocks based on returns across multiple holding periods, outperforming related baselines on relevant metrics.
Machine Learning
@@ 1 Introduction
footnotetext: Department of Computer Science and Automation, Indian Institute of Science, Bengaluru, India - 560012.footnotetext: Correspondence to: Ambedkar Dukkipati $<$ambedkar@iisc.ac.in$>$
Figure 1: Depicts the interdependence of stock trends between stocks in the same class. We observe the normalized stock prices of various stocks during the COVID-19 pandemic from January to May 2020 on the NSE. Specifically, after analyzing assets in the pharmaceutical sector, namely CIPLA, DIVISLAB, and DRREDDY, we observed a notable positive trend following the pandemic event. Conversely, assets in the banking industry, such as PNB, INDUSINDBANK, and ICICIBANK, experienced a negative trend during the same period.
One of the most important forecasting problems is related to stock price movements. Contemporary methods model this as time-series forecasting and emply powerful sequence models with deep neural networks.
Existing techniques, such as graph-based embeddings (Li et al., 2019), graph neural networks (Matsunaga et al., 2019), attention mechanisms (Cheng et al., 2018; Qian et al., 2024), event-driven trading (Zhou et al., 2021) and time-aware embeddings (Feng et al., 2019) effectively model the complex relations between the stocks.
Available forecasting solutions in the literature often rely on historical data patterns without considering the external influences that can change with time. However, it is well known that the pattern of evolution of stock prices is not an isolated time-series but a part of a broader financial and economic environment that encompasses various external events, such as government regulations, an outbreak of war, and diseases. These elements introduce a level of unpredictability that makes forecasting challenging, even if one considers the stock interdependencies. Though event-driven trading and decisions based on sentiment analysis of news are well known, there is a notable lack of research on sequential time-series learning that incorporates such external dynamic events.
@@
To study the problem of forecasting in the presence of external influence, we choose the domain of stock markets and propose a new approach that uses temporal knowledge graphs (KG). Considering that no such KGs are available, we construct relevant knowledge graphs and propose representation learning techniques to set benchmark results. Our constructed temporal KGs integrate diverse components, including first and second-order stock relationships, corporate events, macroeconomic indicators, events derived from financial statements, financial ratios disclosed during earnings calls, and analyst sentiments. Notably, all of these elements are equipped with time-aware attributes. The capacity to incorporate entities of varying categories, such as Person, Stock Assets, Products, and Alliance, results in an intrinsically heterogeneous knowledge graph, with these heterogeneous entities acting as mediators or connectors, enabling the modeling of complex second-order relationships in the graph. This constructed graph, named Stock Market Knowledge Graph ($\mathsf{SmKG}$), aims to encapsulate all the pertinent information related to stock assets within a given share market. We use this knowledge graph to construct an algorithm that models market movements along with the occurrence of external events that affect them, resulting in a holistic approach to modeling markets. As part of our method, we model the relations in the temporal knowledge graph as events of a heterogeneous Hawkes process on the graph to learn useful embeddings that aid stock price analysis.
Our key contributions are as follows.
  1. (i)
@@
  2. (ii)
    For the task of stock prediction, we construct temporal knowledge graphs that capture relevant macroeconomic and financial information related to the NSE and NASDAQ stock markets and use them to enhance sequential forecasting models.
  3. (iii)
    To this end, we propose a time-aware knowledge graph representation technique that leverages a heterogeneous graph convolution operator, along with a temporal point process-based method to model external events. We integrate these embeddings with transformer-based sequential embeddings to obtain accurate rankings of future stock performance over different time horizons.
  4. (iv)
    We empirically demonstrate that our method outperforms related baselines in ranking stocks across three different holding periods. Additionally, we conduct ablation studies and further experiments to analyze the effectiveness of different components of our model.
@@ 2 Related Work
Earlier studies (Zhang et al., 2017) on stock movement patterns employ Recurrent Neural Networks (RNNs), relying solely on historical stock price data for analysis, which is inadequate due to ignoring the influence of numerous uncertain political-economic factors prevalent in the real world. In recent times, advancements in automated trading strategies have emerged, exploring event-driven and news-driven trading approaches by detecting relevant events in news articles (Hu et al., 2018; Zhou et al., 2021) and social media texts (Li et al., 2022b). However, these methods assume that stock movements are independent of one another, restricting the ability to model interactions among different stocks. Further, events extracted from raw texts lack the necessary background knowledge pertaining to the entities and relationships they reference.
A new line of research direction focuses on employing graph-based methodologies to represent pairwise relationships between stocks using metadata such as stock-industry information and various connections between companies, such as supply chain associations and links between CEOs and founders. These methodologies tackle stock movement prediction by formulating it as a node classification problem and enhancing traditional historical price-based prediction models through the utilization of graph-based deep learning approaches. Chen et al. (2018) constructs graphs representing shareholder relationships between stocks and utilizes a graph neural network-based model to learn stock movements. Later works (Feng et al., 2019; Matsunaga et al., 2019; Qian et al., 2024) improved the approach by incorporating additional inter-stock relations, such as sector industry connections, supply chain relationships, and inter-executive relationships. They propose a time-aware graph convolutional network that effectively models these inter-stock relations. Ang & Lim (2022) extended this work to incorporate news text data along with inter-stock relations and train their model to generate forecasts of means and volatilities of stock returns over a horizon. Li et al. (2022a) constructs a heterogeneous multi-modal graph that incorporates historical price data, stock news, and technical indicators to model stock trends. Sawhney et al. (2020) and Sawhney et al. (2021) model higher-order inter-stock relationships by formulating them as hypergraphs. However, despite the efficient modeling of different types of data in isolation, there is currently no existing work that comprehensively captures all of these data modalities in a unified framework. Incorporating external knowledge has been studied in the context of natural language processing. Previous works have tackled the issue of grounding natural language to knowledge by aligning or combining language and knowledge graph representations (Wang et al., 2023).
In our method, we model relations on the external dynamic knowledge graph as events of a Hawkes process on a graph, using the learned embeddings for improved predictions. Neural Temporal Point Processes (Mei & Eisner, 2017) have been used in a wide variety of applications to model temporal event streams and predict the occurrence and features of future events. Such processes have further been used to model dynamic interactions between different entities, leading to applications in graph edge prediction and relation forecasting in dynamic graphs (Gracious et al., 2025; Trivedi et al., 2019; Ji et al., 2021). Edge prediction has also been studied without considering them as temporal processes using latent space models (Gracious et al., 2021) and attention mechanisms (Xu et al., 2020).
@@
Figure 2: Different types of relations: (left) without timestamp, (center) with timestamp, and (right) with time interval.
@@ 3 Temporal Knowledge Graph Construction
The construction of the knowledge graphs $\mathsf{SmKG}$ involves data acquired from various sources pertaining to the NASDAQ and NSE share markets, mentioned below. Specifically, the stocks listed under the NASDAQ100, S&P500, and NIFTY500 indices are considered. The knowledge graphs are formed to encompass a time span of 20 years, commencing from January 01, 2003, and concluding on December 30, 2022. Statistics of the knowledge graphs for NASDAQ and NSE share markets are shown in Table 1.
Table 1: Statistics of the NASDAQ and NSE $\mathsf{SmKG}$s
$\mathsf{SmKg}$ NASDAQ NSE
@@
Figure 3: Visualization of a subgraph from the constructed $\mathsf{SmKG}$s for (left) NASDAQ and (right) NSE share markets. The entities and relations are highlighted based on their types.
General News News articles were acquired by scraping data from two websites, namely “Seeking Alpha” for NASDAQ and “Money Control” for NSE, using the Selenium library in Python [^1] The news articles are obtained by scraping the webpages https://seekingalpha.com/symbol/{stock_ticker}/news for NASDAQ and https://www.moneycontrol.com/company-article/{stock_name}/news/{stock_id} for NSE. A total of 220,645 news articles were collected for NASDAQ and 272,808 for NSE. These articles comprise several components, including the headline, body text, publication time, associated stock asset, related sector, and additional taglines, and contain various types of information such as corporate events, earnings reports, analyst and investor sentiments, etc. To identify the events mentioned in the news articles, a rule-based pattern recognition system was employed for a predefined set of events. The identified events are then added to the constructed knowledge graph $\mathsf{SmKG}$ in the form of quadruple $(h,r,t,[\tau_{s}])$, where $\tau_{s}$ is the news article publication time.
Financial Statements and Ratios The financial statements of companies listed under NASDAQ and NSE were extracted by scraping data from the websites “Macrotrends” and “Top Stock Research” using the Selenium library in Python [^2] The financial statements are obtained from www.macrotrends.net and https://www.topstockresearch.com/ for NASDAQ and NSE respectively.. The collected financial data for each stock consists of various essential components, including the balance sheet, income statement, cash flow statement, and financial ratios. This data is acquired at both annual and quarterly frequencies. Events are derived from the financial statements to identify financial trends by examining whether specific financial parameters exhibit an upward or downward trend. These event trends are then incorporated into the knowledge graph as quintuples, where the interval is defined by the beginning and end of the corresponding financial year or quarter.
Macro-Economic Indicators Macroeconomic indicators provide a snapshot of the overall economic health of a country, encompassing factors such as GDP growth, inflation rate, unemployment rate, trade balance, and fiscal deficit. This information is obtained from “Macrotrends” for both the US and India in which the respective share markets NASDAQ and NSE are located [^3] The macro-economic indicators are obtained from www.macrotrends.net.. The trends in the indicators are identified as events with intervals defined by the beginning and end of the reported year.
Corporate Events Corporate events such as dividend announcements and stock splits are acquired from “Yahoo Finance” for NASDAQ and NSE. These events are represented within the knowledge graph as quadruples, with the timestamp being the date on which the event occurred.
Inter-stock Relationship The inter-stock relationships can be considered to be of two types: first-order and second-order relationships. If there exists a statement with company $i$ as the subject and company $j$ as the object, then they have a first-order relation. Furthermore, companies $i$ and $j$ have a second-order relation if they have statements that share the same object (Feng et al., 2019). For instance, “Apple” and “Foxconn” are in different stages in the production of the product “iPhone”. This data was extracted from “Wikibase,” where the knowledge from “Wikipedia” is stored in the form of triples. A preselected set of relations is identified consisting of $6$ types of first-order relations and $52$ types of second-order relations, which are then filtered from Wikibase.
Sector Classification Each stock asset is classified into a four-level hierarchical system: Sector, Industry group, Industry, and Sub-Industry by Global Industry Classification Standard (GICS) for NASDAQ and Macroeconomic Sector, Sector, Industry, and Basic Industry by NSE Indices for NSE. These relations are added as triples with the asset as head, class as tail, and classification level as relation to the constructed knowledge graph.
@@ 4 Proposed Learning Mechanism
Figure 4: TA-HKGE model architecture
4.1 Overview
The input to the model $X^{T}$ at time $T$ consists of $X^{T}=\{\mathcal{P}_{i}^{\tau\in(T-W,T]},G_{\tau_{0}},G_{T}\}$, where $\mathcal{P}_{i}^{\tau\in(T-W,T]}$ is the window of historical stock price data for $W$ previous days of stock $i$ at day $T$, $G_{\tau_{0}}$ is the static subgraph of $\mathsf{SmKG}$, $G_{\tau_{T}}$ is the temporal positive subgraph of $\mathsf{SmKG}$ at time $T$, and $N$ is the total number of stocks considered in the dataset.
The objective is to learn an encoder $\mathcal{E}(X^{T})$ to obtain output embedding $\mathcal{H}^{T}=\left(H_{i}^{T}\right)_{1\leq i\leq N}$. We construct his encoder using multiple components: a Sequential Encoder $\mathcal{E}_{s}(\mathcal{P}_{i}^{\tau\in(T-W,T]})=E_{s}$ of the stock prices, a set of temporal process embeddings $e_{\mathrm{tpp}}$ that are learned to optimally model the events as a heterogeneous Hawkes process, and a Relational Encoder $\mathcal{E}_{r}(G_{\tau_{0}},G_{T},E_{s},e_{\textrm{tpp}})=E_{h}$. These are then passed through the prediction layer $\mathbf{Y}^{\Delta}_{\mathbf{W}_{f}}$ to obtain a vector of ranking scores for the stocks at time step $T+\Delta$.

### v2

# Predictive AI with External Knowledge Infusion: Datasets and Benchmarks for Stock Markets

@@ Abstract
Fluctuations in stock prices are influenced by a complex interplay of factors that go beyond mere historical data. These factors, themselves influenced by external forces, encompass inter-stock dynamics, broader economic factors, various government policy decisions, outbreaks of wars, etc. Furthermore, all of these factors are dynamic and exhibit changes over time. In this paper, for the first time, we tackle the forecasting problem under external influence by proposing learning mechanisms that not only learn from historical trends but also incorporate external knowledge from temporal knowledge graphs. Since there are no such datasets or temporal knowledge graphs available, we study this problem with stock market data, and we construct comprehensive temporal knowledge graph datasets. In our proposed approach, we model relations on external temporal knowledge graphs as events of a Hawkes process on graphs. With extensive experiments, we show that learned dynamic representations effectively rank stocks based on returns across multiple holding periods, outperforming related baselines on relevant metrics.
@@ 1 Introduction
The development of learning algorithms for forecasting events is an important component in building AI systems with a wide range of applications, including in financial markets and healthcare.
Figure 1: Interdependence of stock trends between stocks in the same class during COVID-19 (January–May 2020, NSE). Pharmaceutical stocks (CIPLA, DIVISLAB, DRREDDY) trended positively, whereas banking stocks (PNB, INDUSINDBANK, ICICIBANK) declined.
Recent progress in these areas has been notable, and methods involving a combination of power sequence models, such as transformers, and elegant probabilistic models, such as temporal point processes, have made enormous progress (Mei and Eisner, 2017; Gracious et al., 2021; Gracious and Dukkipati, 2023; Gracious et al., 2025; Gracious and Dukkipati, 2025). However, nearly all existing methods fall short of adequately accounting for or integrating time-varying external factors that influence the outcomes. For instance, stock price variations are affected by a complex interplay of factors that extend beyond mere historical patterns of stock prices. These elements can be shaped by external influences, such as prevailing economic conditions, shifts in governmental policies, and incidents of conflict. Additionally, these factors can be dynamic, with some changing rapidly, others very slowly, some very frequently, and others infrequently.
@@
To study this problem, we develop an approach that captures external influences via temporal Knowledge Graphs (KG) and learns historical patterns via temporal point processes. Considering that such specific temporal KGs are not available in the literature, we develop temporal KGs for financial markets for the machine learning community and propose representation learning techniques to establish benchmark results. The temporal KGs we constructed integrated diverse components, including first- and second-order stock relationships, corporate events, macroeconomic indicators, events derived from financial statements, financial ratios disclosed during earnings calls, and analyst sentiments. Notably, all these elements are equipped with time-aware attributes. The capacity to incorporate entities of varying categories, such as Person, Stock Assets, Products, and Alliance, results in an intrinsically heterogeneous knowledge graph. We refer to these KGs as $\mathsf{STOCKnowledge}$ aims to encapsulate all pertinent information related to stock assets within a given share market.
Contributions
@@
(2) For the task of stock prediction, we construct temporal knowledge graphs that capture relevant macroeconomic and financial information related to the NSE and NASDAQ stock markets and use them to enhance sequential forecasting models.
(3) To this end, we propose a time-aware knowledge graph representation technique that leverages a heterogeneous graph convolution operator, along with a temporal point process-based method to model external events. We integrate these embeddings with transformer-based sequential embeddings to obtain accurate rankings of future stock performance over different time horizons.
(4) We empirically demonstrate that our method outperforms related baselines in ranking stocks across three different holding periods. Additionally, we conduct ablation studies and further experiments to analyze the effectiveness of the different components of our model.
@@ 2 Related work
Forecasting has traditionally used time-series models that predict future values solely from historical numerical data, assuming that the target variable evolves as a function of its own past. Deep learning techniques using models such as RNNs (Hewamalage et al., 2021), LSTMs (Elsworth and Güttel, 2020; Kong et al., 2025), and transformer-based architectures (Zhou et al., 2021; Liu et al., 2024) enable the capture of long-range dependencies and complex non-linear patterns that classical models struggle to represent. However, these approaches often assume that all relevant signals are contained within the target series and a limited set of covariates, overlooking the broader context in which the data arise. This limits their ability to handle situations in which external events (e.g., pandemics, geopolitical shifts) or inter-entity interactions (e.g., firm-to-firm relationships) cause abrupt or correlated changes. While recent work has begun to incorporate covariates, there remains a gap in systematically modeling rich, temporally dynamic external influences.
To address this gap, recent studies have incorporated unstructured data sources into forecasting models. These include news articles (Wang et al., 2024a), social media posts (Karlemstrand and Leckström, 2021), financial reports (Zheng et al., 2019), and product reviews (Liu et al., 2023), which offer valuable contextual signals that complement numerical time series. Text-Guided Time Series Forecasting (TGTSF)(Xu et al., 2024) uses textual cues, such as channel metadata or news content, to guide predictions. Con (Chattopadhyay et al., 2025) combines temporal and textual modalities using cross-attention to achieve improved results. In a multimodal direction, MM-Forecast (Li et al., 2024) integrates images and other modalities for temporal-event prediction. These approaches typically treat unstructured signals as flat texts, failing to capture the relationships between entities or how they evolve over time. This limits their ability to reason about structured causal chains or model how influences unfold temporally.
Incorporating structural knowledge into learning problems has been studied for natural language problems (Annervaz et al., 2018). Recently, this technique has been applied to forecasting problems. This includes time-varying categorical variables, continuous external signals, and relational knowledge captured using graphs and knowledge graphs. For example, TimeXer (Wang et al., 2024b) enhances transformers with patch-wise and variate-wise attention to handle structured exogenous inputs in multivariate time-series. ChronosX (Arango et al., 2025) fuses historical and future auxiliary inputs using dedicated networks to adapt pretrained models without retraining the core models. Beyond tabular features, recent studies have constructed graphs (Cini et al., 2023) and knowledge graphs (Jain et al., 2023) to model prior knowledge and event dynamics that influence targets. This structured context supports better generalization across domains and improves interpretability by linking the forecasts to causally relevant external signals. However, these efforts are often static or limited to coarse-grained relationships and do not explicitly model the evolving temporal interactions between entities.
Our work addresses this gap by constructing temporal knowledge graphs that encode fine-grained, time-indexed external events and using point-process modeling to dynamically represent their influence on forecasting targets.
3 Problem formulation
Consider a forecasting problem on a multivariate time series
$\mathcal{P}=\left(\mathcal{P}_{i}^{\tau}:i=1,\dots,N\right)_{\tau},$
with $N$ values at each time $\tau$ in some discrete sequence of time points $\tau_{1},\tau_{2}\dots$. At time $T$, using the history of the series till that time, we wish to forecast
$Y^{T+\Delta}=\displaystyle\mathbf{Y}^{\Delta}\left(\left(\mathcal{P}_{i}^{\tau}\right)_{1\leq i\leq N}^{\tau_{1}<\tau\leq T}\right),$
where $Y^{T+\Delta}$ is some property of the $N$ values $\left(\mathcal{P}_{i}^{T+\Delta}\right)_{1\leq i\leq N}$ of the time series that occur at time $T+\Delta$. For the stock forecasting problem, $\mathcal{P}_{i}^{\tau}$ denotes the price of stock asset $i$ at time $\tau$, and $Y^{T+\Delta}$ is the ranking of the stock prices of all assets at time $T+\Delta$.
As mentioned earlier, we model the temporal knowledge of external events that can influence this series as a temporal KG. For this, we consider the elements $\left(\mathcal{P}_{i}^{\tau}\right)_{\tau}$ of this time series to occur on a set of $N$ nodes of a temporal KG $G$. While $(\mathcal{P}_{i}^{\tau})$ represents a multivariate series that evolves continuously at every time step, the knowledge graph $G$ has a static component that contains information about all the fixed relations between entities in the graph and a dynamic component that captures the occurrence of events across different time scales. Each event is represented as an edge between the nodes of the graph, valid for a certain time period based on the nature of the event.
To this end, we extended the standard static knowledge graph, consisting of triples of the form $(h,r,t)$, to a temporal knowledge graph by adding a time dimension for the relations, resulting in quintuples of the form $(h,r,t,[\tau_{s},\tau_{e}])$, where $h$ and $t$ are the head and tail entities, respectively, $r$ is the type of relationship between them, and $\tau_{s},\tau_{e}]$ is the time interval in which the triple remains valid. If the triple is valid for only the time step $\tau$, then it is represented as a quadruple $(h,r,t,[\tau])$.
Thus, the temporal knowledge graph $G$ is a union of subgraphs
$G=G_{\tau_{0}}\cup\bigcup_{j=1}^{T}G_{\tau_{j}},$
where $G_{\tau_{0}}$ is the static subgraph and $G_{\tau_{j}}$ is the set of relations valid at time $\tau_{j}$. Therefore, the input $X^{T}$ to our model at time $T$ is of the form
$X^{T}=\left(\mathcal{P}_{1\leq i\leq N}^{T-W<\tau\leq T},G_{\tau_{0}},G_{T}\right),$
where $\mathcal{P}_{i}^{T-W\leq\tau\leq T}$ is the window of historical stock price data for $W$ previous days of stock $i$ at day $T$, $G_{\tau_{0}}$ is the non-temporal positive subgraph of $G$, and $G_{\tau_{j}}$ is the temporal positive subgraph of $G$ at time $\tau_{j}$.
The objective is to learn embeddings $\mathcal{E}(X^{T})$ that can be used to compute scores
$\widehat{Y}^{T+\Delta}=\mathbf{Y}^{\Delta}_{\mathbf{W}_{f}}(\mathcal{E}(X^{T}))$
to rank the best-performing stocks at future time $T+\Delta$. The embedding function $\mathcal{E}$ along with the final prediction layer parameters $\mathbf{W}_{f}$ are learned using an objective that is a combination of knowledge graph embedding consistency loss and ranking losses that align the output with the ground truth values $\mathcal{P}_{i}^{T+\Delta}$. The components of this loss function, as well as the detailed equations, are described in Section 5.5 and the Appendix A. The learned model is used to predict the top $1$ and $5$ best-performing stocks $\Delta$ time steps in the future. The quality of these predictions is quantified using various relevant metrics, such as return ratios, Sharpe ratios, and accuracy, as described in Appendix B.
@@ 3 Problem formulation
Figure 2: Different types of relations: (left) without a timestamp, (center) with a timestamp, and (right) with a time interval.
@@ 4 Temporal Knowledge Graph construction
The construction of the KGs $\mathsf{STOCKnowledge}$ involved data acquired from various sources pertaining to the NASDAQ and NSE share markets. Temporal KGs are constructed for stocks listed in NASDAQ100, S&P500, and NIFTY500, spanning a period of 20 years, commencing from January 01, 2003, and concluding on December 30, 2022.
General News News articles were acquired by scraping data from two websites, namely “Seeking Alpha” for NASDAQ and “Money Control” for NSE, using the Selenium library in Python [^1] The news articles are obtained by scraping the webpages https://seekingalpha.com/symbol/{stock_ticker}/news for NASDAQ and https://www.moneycontrol.com/company-article/{stock_name}/news/{stock_id}. A total of 220,645 news articles were collected for the NASDAQ and 272,808 for the NSE. These articles comprise several components, including the headline, body text, publication time, associated stock asset, related sector, and additional taglines, and contain various types of information, such as corporate events, earnings reports, and analyst and investor sentiments. A rule-based pattern-matching system was used to identify events mentioned in the articles using a predefined set of events, which were then added to the knowledge graph.
Figure 3: TA-HKGE model architecture
Financial Statements and Ratios The financial statements of companies listed on NASDAQ and NSE were extracted by scraping data from websites “Macrotrends” and “Top Stock Research” using Selenium [^2] The financial statements are obtained from www.macrotrends.net and https://www.topstockresearch.com/ for NASDAQ and NSE, respectively.. This data includes essential components for each stock, including the balance sheet, income statement, cash flow statement, and other financial ratios. Events are derived from financial statements to identify the annual and quarterly trends of these financial parameters. These trends were then incorporated into the knowledge graph as quintuple events, where the interval was defined by the beginning and end of the corresponding financial year or quarter.
Macro-Economic Indicators Macroeconomic indicators provide a snapshot of the general economic health of a country, including factors such as GDP growth, inflation rate, unemployment rate, trade balance, and fiscal deficit. This information is obtained from “Macrotrends” for both the US and India, in which the respective share markets NASDAQ and NSE are located [^3] The macro-economic indicators are obtained from www.macrotrends.net.. Trends in indicators were identified as events with intervals defined by the beginning and end of the reported year.
Corporate Events Corporate events, such as dividend announcements and stock splits, are sourced from Yahoo Finance for the NASDAQ and NSE.
Table 1: Statistics of the NASDAQ and NSE $\mathsf{STOCKnowledge}$s
$\mathsf{STOCKnowledge}$ NASDAQ NSE
@@
Inter-stock Relationship Inter-stock relationships can be classified into two types: first-order and second-order relationships. If there is a statement with company $i$ as the subject and company $j$ as the object, then they have a first-order relationship. Furthermore, companies $i$ and $j$ have a second-order relationship if they have statements that share the same object (Feng et al., 2019). For instance, “Apple” and “Foxconn” are in different stages in the production of the product “iPhone”. This data was extracted from “Wikibase,” where the knowledge from “Wikipedia” is stored in the form of triples. A preselected set of relations was identified, consisting of $6$ types of first-order relations and $52$ types of second-order relations, which were then filtered from Wikibase.
Sector Classification Each stock asset is classified into a four-level hierarchical system: Sector, Industry group, Industry, and Sub-Industry by Global Industry Classification Standard (GICS) for NASDAQ, and Macroeconomic Sector, Sector, Industry, and Basic Industry by NSE Indices for NSE. These relations are added as triples with the asset as the head, class as the tail, and classification level as the relation to the constructed knowledge graph.
@@ 5 Proposed Learning Mechanism
The input to the model is $X^{T}$ at time $T$ consists of
$X^{T}=\{\mathcal{P}_{i}^{\tau\in(T-W,T]},G_{\tau_{0}},G_{T}\},$
where $\mathcal{P}_{i}^{\tau\in(T-W,T]}$ is the window of historical stock price data for $W$ previous days of stock $i$ at day $T$, $G_{\tau_{0}}$ is the static subgraph of $\mathsf{STOCKnowledge}$, $G_{\tau_{T}}$ is the temporal positive subgraph of $\mathsf{STOCKnowledge}$ at time $T$, and $N$ is the total number of stocks in the dataset.
The objective is to learn an encoder $\mathcal{E}(X^{T})$ to obtain output embeddings that can be used to predict the ranking of various stocks at future time steps. We construct this encoder using multiple components: a Sequential Encoder
$\mathcal{E}_{s}(\mathcal{P}_{i}^{\tau\in(T-W,T]})=E_{s}$
of the stock prices, a set of temporal process embeddings $e_{\mathrm{tpp}}$ that are learned to optimally model the events as a heterogeneous Hawkes process, and a Relational Encoder
$\mathcal{E}_{r}(G_{\tau_{0}},G_{T},E_{s},e_{\textrm{tpp}})=E_{h}.$
These are then passed through a prediction layer $\mathbf{Y}^{\Delta}_{\mathbf{W}_{f}}$ to obtain a vector of ranking scores for the stocks at time step $T+\Delta$.